Information from the financial statements of Ames Fabricators, Inc., included the following: December 31 2018 2017 Common shaNumerator / Denominator Earnings per Share / Basic EPS Diluted EPS / II
Answer #1

Basic EPS = (Net Income - Preferred dividend) / No. of common stock

2018 = (600000 - 18000*5) / 100000 = $5.10 per share

Diluted EPS = (Net Income + Interest saving net of tax) / (Weighted average outstanding common share + potential common shares)

= (600000 + 1000000*10%*60%) / (100000+42000+20000)

= $4.07 per share